# Aligning Individual and Collective Objectives in Multi-Agent Cooperation

### v1

@@ 4.1 Differentiable Mixed-motive Game
We first formulate the mixed-motive game as a differentiable game. Specifically, the differentiable mixed-motive game (DMG) is defined as a tuple $({\mathcal{N}},{\bm{w}},{\bm{\ell}})$, where $\ell_{i}\in{\bm{\ell}}$ is at least twice the differentiable loss function for the agent $i$. Differentiable losses exhibit the mixed motivation property: minimization of individual losses can result in a conflict between individuals or between individual and collective objectives (e.g., maximizing social welfare or beating opponents in basketball matches). The primary objective of DMG is to optimize the collective objective by minimizing collective loss, denoted as $\ell_{c}$, while simultaneously taking into account the need to minimize individual objectives, represented by ${\bm{\ell}}$. Before delving into our interest alignment methods, let us first introduce some notations. We use ${\bm{\xi}}_{c}({\bm{w}})$ to refer the gradient of collective loss: ${\bm{\xi}}_{c}({\bm{w}})=\left(\nabla_{\mathbf{w}_{1}}\ell_{c},\ldots,\nabla_{\mathbf{w}_{n}}\ell_{c}\right).$
Nonetheless, direct optimization of individual or collective losses can lead to convergence issues and a mismatch between individual and collective interests. It is improbable that by minimizing only the individual loss of each agent, a collective minimum will be achieved (Leibo et al., 2017; McKee et al., 2020b). On the other hand, optimizing collective loss might produce improved collective outcomes, but at the risk of overlooking individual interests. This phenomenon is also demonstrated in Example 4.1. Furthermore, the local convergence of the gradient descent on singular collective functions is not always guaranteed (Balduzzi et al., 2018). This also typifies the problem of credit assignment in MARL.
Alignment Dilemma in DMG. Here, we provide an example of a two-player differentiable mixed-motive game and the optimization results of the aforementioned methods.
@@
We provide a visual representation of optimization trajectories using a series of methods to investigate the learning dynamics involved in resolving the toy two-player differentiable mixed-motive game. In particular, we first implement simultaneous optimization of individual losses (Simul-Ind) with respect to each parameter, defined by the learning rule ${\bm{w}}_{i}={\bm{w}}_{i}-\gamma{\bm{\xi}}_{i}$, for $i\in\{1,2\}$, where ${\bm{\xi}}_{i}=\nabla_{{\bm{w}}_{i}}\ell_{i}$. Simultaneous optimization of collective loss (Simul-Co) replaces the individual gradient with the collective gradient ${\bm{\xi}}_{c}$ of collective loss $\ell_{c}=\ell_{1}+\ell_{2}$. Furthermore, in order to investigate the learning dynamics of prevalent gradient modification optimization approaches such as SGA and CGA for the two-player differentiable mixed-motive game, we modify the learning rule by integrating the respective adjusted gradient as delineated in Section 3.
@@ 4.2 Altruistic Gradient Adjustment
In this section, we propose the alignment of individual and collective interests towards achieving stable fixed points of collective objectives within the context of mixed-motive cooperation. Revisit that the learning dynamics of simultaneous optimization on collective loss is given by ${\bm{w}}={\bm{w}}-\gamma{\bm{\xi}}_{c},$ where ${\bm{\xi}}_{c}$ represents the gradients of collective loss $\ell_{c}$ respectively, with respect to the parameters of each player ${\bm{w}}_{i}\in{\bm{w}}$. Although Simul-Co efficiently enhances collective performance, its shortcomings lie in not guaranteeing stable fixed points (Balduzzi et al., 2018) and overlooking individual interests as discussed in Example 4.1.
To tackle this challenge, we introduce the Altruistic Gradient Adjustment (AgA) method with the goal of aligning individual and collective objectives while seeking stable fixed points. Unlike existing gradient adjustment methods (Mescheder et al., 2017; Balduzzi et al., 2018; Letcher et al., 2019; Chen et al., 2023) that primarily focus on identifying the stable fixed points of individual losses, our proposed AgA method integrates the individual gradient into the collective gradient and ensures convergence to stable fixed points of the collective loss. Specifically, AgA is defined as follows.
@@
Figure 2: Illustration of Corollary 4.3. In case 1, within an unstable fixed point’s neighborhoods, an appropriate selection of the $\lambda$ sign push AgA to evade the unstable fixed point and pull towards a stable fixed point in its neighborhoods as shown in case 2.
Note that the Hessian matrix ${\bm{H}}_{c}$ is a symmetric matrix. To simplify the notation, we use $\nabla{\mathcal{H}}_{c}$ to denote the term $\nabla_{\bm{w}}\left(\frac{1}{2}\|{\bm{\xi}}_{c}\|^{2}\right)$, i.e., ${\bm{H}}_{c}^{T}{\bm{\xi}}_{c}$.
Intuitively, AgA is designed to involve both collective and individual objectives, in order to harmonize the interests of individuals and the team. The alignment parameter $\lambda$ adjusts the balance between the individual and the collective objective. Furthermore, we minimize ${\mathcal{H}}_{c}=\frac{1}{2}\|{\bm{\xi}}_{c}\|^{2}$ to converge to a fixed point. This objective is reflected in the gradient term ${\bm{H}}_{c}^{T}{\bm{\xi}}_{c}$ in $\tilde{{\bm{\xi}}}$. Assuming that ${\bm{H}}_{c}$ is invertible, we can deduce that $\nabla{\mathcal{H}}_{c}={\bm{H}}_{c}^{T}{\bm{\xi}}_{c}=0$, if and only if ${\bm{\xi}}_{c}=0$. Therefore, minimizing ${\mathcal{H}}_{c}$ will guarantee the convergence to a fixed point (Letcher et al., 2019).
Impact of the Sign of $\lambda$. While the modified AgA gradient introduces additional complexity, we theoretically demonstrate that an appropriate choice of the sign of $\lambda$ ensures that the AgA gradient $\tilde{{\bm{\xi}}}$ is attracted towards a stable fixed point. In contrast, when dealing with an unstable fixed point, AgA is pushed away the point.
Before we delve into the corollary of the sign selection for $\lambda$, we first establish some foundational definitions. The inner product of vectors ${\bm{a}}$ and ${\bm{b}}$ are denoted by $\langle{\bm{a}},{\bm{b}}\rangle$. If the Hessian matrix ${\bm{H}}$ is non-negative-definite, $\langle{\bm{\xi}}_{c},\nabla{\mathcal{H}}\rangle\geq 0$ for a non-zero ${\bm{\xi}}_{c}$. Analogously, if ${\bm{H}}$ is negative-definite, $\langle{\bm{\xi}}_{c},\nabla{\mathcal{H}}\rangle<0$ for a non-zero ${\bm{\xi}}_{c}$. Lastly, the angle between the two vectors ${\bm{a}}$ and ${\bm{b}}$ is denoted by $\theta({\bm{a}},{\bm{b}})$.
The following corollary states that AgA with a suitable sign of $\lambda$ will bring the adjusted gradient $\tilde{{\bm{\xi}}}$ point closer to a stable fixed point and escape from unstable fixed points.
@@
If we let the sign of $\lambda$ satisfy $\lambda\cdot\langle{\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c}\rangle(\langle{\bm{\xi}},\nabla{\mathcal{H}}_{c}\rangle+\|\nabla{\mathcal{H}}_{c}\|^{2})\geq 0$, the optimization process using the AgA method exhibit following attributes. Primarily, in neighborhood of fixed point of collective objective, if the point is stable (${\bm{H}}_{c}\succeq 0$), the AgA gradient will be pulled toward this point, which means that $\theta(\tilde{{\bm{\xi}}},\nabla{\mathcal{H}}_{c})\leq\theta({\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c})$. On the other hand, if the point represents unstable equilibria where ${\bm{H}}_{c}\prec 0$, the AgA gradient will be pushed out of the point, indicating that $\theta(\tilde{{\bm{\xi}}},\nabla{\mathcal{H}}_{c})\geq\theta({\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c})$.
@@
The proof is provided in Appendix A. Fig. 2 presents a illustrative visualization of Corollary 4.3. The figure depicts two scenarios: the neighborhoods of stable and unstable fixed point (denoted by a star), respectively. In the neighborhood of an unstable fixed point of the collective objective, as depicted in Case 1, our AaA gradient $\tilde{{\bm{\xi}}}$ is pushed out of the region, resulting in a faster escape compared to the original collective gradient ${\bm{\xi}}_{c}$. This behavior is exemplified by the green arrow. Conversely, Case 2 illustrates the scenario of a stable fixed point, wherein our AaA gradient $\tilde{{\bm{\xi}}}$ is pulled towards the stable equilibrium, exemplified by the red arrow.
Algorithm 1 Altruistic Gradient Adjustment (AgA)
3:  ${\bm{\xi}}_{c}\leftarrow[grad(\ell_{c},{\bm{w}}_{i})\ for\ {\bm{w}}_{i}\in{\bm{w}}]$
Implementation of MARL Algorithms with AgA. The concrete implementation of altruistic gradient adjustment is elaborated in Algorithm 1. AgA can be easily incorporated into any centralized training decentralized execution (CTDE) framework of MARL. In a typical CTDE framework, each agent, denoted as $i$, strives to learn a policy that employs local observations to prompt a distribution over personal actions. The unique characteristic during the centralized learning phase is that agents gain supplementary information from their peers that is inaccessible during the execution stage. Furthermore, a centralized critic often operates in tandem in CTDE settings, leveraging shared information and evaluating individual agent policies’ potency. To incorporate our proposed AgA algorithm into CTDE frameworks, supplementary reward exchanges between agents are required, offering the essential inputs for Algorithm 1. The individual losses are calculated from individual rewards conforming to established MARL algorithm framework, such as IPPO (de Witt et al., 2020) or MAPPO (Yu et al., 2022). Here, the collective loss $\ell_{c}$ is calculated according to individual rewards; for example, it could be a sum of individual rewards or an elementary transformation thereof. Furthermore, the AgA algorithm requires a magnitude value of the alignment parameter $\lambda$, which should be a positive entity. The sign of $\lambda$ is then calculated in line 5 of Algorithm 1 and is based on each gradient component calculated in lines 2-4. Subsequently, the algorithm produces the adjusted gradient $\tilde{{\bm{\xi}}}$. $\tilde{{\bm{\xi}}}$ is then distributed across each corresponding parameter, and optimization is carried out using a suitable optimizer such as Adam optimizer (Kingma & Ba, 2015).
(a) Reward Contour of Player 1
(b) Reward Contour of Player 2
Figure 3: The comparison between AgA (shown in red) and AgA without sign alignment (AgA w/o Sign, in purple) trajectories spans 40 steps, marked at every tenth step. Norm gradients are represented with blue arrows. Starting from the 14th step, sign alignment pull the update direction. Consequently, AgA is about 6 steps ahead of AgA w/o Sign at the trajectory’s end.
4.3 Alignment Efficiency of AgA: A Toy Experiment
@@
To address the two-player DMG as outlined in Example 4.1, we incorporate the AgA method into the fundamental gradient descent algorithm. The findings reveal that the AgA method is effective in facilitating the alignment of objectives. The optimization trajectories in both the collective reward landscape and the individual player reward contour are represented in Fig. 0(a), Fig. 0(b), and Fig. 0(c), with brighter colors reflecting higher collective rewards. In contrast to the objective misalignment exhibited by Simul-Co, AgA successfully aligns the agents’ interests, as evidenced in Fig. 0(b) and Fig 0(c). Fig. 0(b) depicts the trajectory of AgA meticulously carving its course along the summit of the individual reward contour. Furthermore, AgA is also shown to improve the reward of Player 2 (as shown in Fig. 0(c)), despite a convergence rate slower than that of Simul-Co. Evidently, the trajectory of AgA (as depicted in Fig. 0(b)) takes 100 steps, whereas Simul-Co takes 65 steps.
@@
In addition, Fig. 3 presents a critical comparison between the trajectories of AgA (in red) and AgA without sign alignment (AgA w/o Sign, depicted in purple), as outlined in Corollary 4.3. This side-by-side comparison covers 40 steps and features highlighted points every tenth step. It provides a visual illustration of the efficiency introduced by sign alignment starting from the 14th step in the trajectories of AgA and AgA without sign alignment. Remarkably, norm gradients are represented by blue arrows, indicating the direction of the fastest updates. As depicted in the figure, AgA with sign selection is aligned towards the fastest update direction, resulting in AgA progressing approximately 6 steps ahead of AgA without Sign at the end of the trajectory.
@@ 5 Experiments
In this section, we conduct a series of experiments to verify the AgA algorithm. First, we provide an overview of the experimental settings and clarify the evaluation metrics employed.
Baselines and Evaluation Metrics. A comparison of experiments involves specific baselines such as simultaneous optimization with individual and collective losses (abbreviated as Simul-Ind and Siml-Co), CGA (Mescheder et al., 2017), SGA (Balduzzi et al., 2018) and SVO (McKee et al., 2020a). In the two-player public goods game experiment, these techniques are simply integrated into the basic gradient ascent algorithm, whereas, for more intricate settings like Cleanup, Harvest (Leibo et al., 2017), and SMAC (Samvelyan et al., 2019), these baselines are infused within various RL algorithms. Detailed operations for implementing these RL algorithms are further elucidated in their respective sections.
We assess collaborative performance in mixed-motive games using a three-tier approach. Primarily, we employ social welfare to gauge overall team performance. Next, an equality metric is introduced to scrutinize fairness in environments such as Cleanup and Harvest. Lastly, the probability of victory against opposition in an SMAC framework determines the probability of the team’s success. Social Welfare is a commonly accepted metric for evaluating the performance of mixed motivation cooperation, representing the total episodic rewards earned by all agents as $SW=\sum_{i=1}^{n}r_{i}$. Social welfare lacks the ability to analyze teamwork in mixed cooperative-competitive contexts. Here, we use the average Win Rate against the built-in computer AI bot in SMAC to assess algorithmic performance. We also involve the Equality due to the measures like social welfare and win rate are limited to a general performance evaluation, ignoring individual interest. Therefore, the Gini coefficient, originally intended as an income equality measure (David, 1968), has been modified to assess reward equality in cooperative AI settings. This study utilizes an expedited Gini coefficient calculation method, with an associated equality metric $E$, formulated as $E:=1-G$. Each individual’s rewards are arranged in ascending order to create a new payoff vector ${\bm{p}}$, and the Gini coefficient is computed as $G=\frac{2}{n^{2}\bar{{\bm{p}}}}\sum_{i=1}^{n}i\left(p_{i}-\bar{{\bm{p}}}\right)$, followed by $E=1-G$. Here, $\bar{{\bm{p}}}$ denotes the mean of the ranked payoff vector ${\bm{p}}$ and $n$ signifies the total number of players. A larger $E$ value implies greater equality.
(a) Cleanup: AgA with different $\lambda$ and AgA/Sign
(b) Cleanup: comparison between AgA and baselines
(c) Harvest: AgA with different $\lambda$ and AgA/Sign
(d) Cleanup: comparison between AgA and baselines
Figure 4: Comparison of Episodic Social Welfare: our proposed AgA methods along with baselines on Cleanup (figure a and b) and Harvest (figure c and d) tasks. Figure (a) and (c) provide an illustration of varying alignment parameter $\lambda$ values in AgA, as well as AgA sans sign selection (AgA/Sign), as applied to Cleanup and Harvest tasks respectively. Subfigures (b) and (d) exhibit the social welfare obtained by AgA utilizing the optimal $\lambda$ value, in comparison to those secured by Simul-Ind, Simal-Co, CGA, SGA and SVO. Bold lines represent the average of the social welfare across three seeds, with the encompassing shaded areas denoting the standard deviation.
Metrics Algorithms
Simul-Ind CGA SGA SVO Simul-Co AgA
${r_{1}}$ 1.13 1.16 1.18 1.10 1.43 1.46
${r_{2}}$ 1.18 1.15 1.14 1.06 1.38 1.45
${SW}$ 2.32 2.31 2.31 2.16 2.81 2.90
${E}$ 0.92 0.93 0.92 0.93 0.94 0.96
Table 1: The comparison of the average individual rewards (denoted as $r_{1},r_{2}$), social welfare (denoted as $SW$) and equality metric (denoted as $E$). Each value is the mean of 50 random runs.

### v2

@@ 4 Method
In this section, we first define differentiable mixed-motive games (DMG) and analyzes the alignment dilemma faced by existing methods, which is described in Section 4.1. We then propose the Altruistic Gradient Adjustment (AgA) algorithm in Section 4.2 to align the individual and collective objectives by modifying the gradient. Finally, a case study demonstrating AgA’s effectiveness in a toy two-player DMG is presented in Section 4.3.
@@ 4.1 Differentiable Mixed-motive Game
We first formulate the mixed-motive game as a differentiable game. Specifically, the differentiable mixed-motive game (DMG) is defined as a tuple $({\mathcal{N}},{\bm{w}},{\bm{\ell}})$, where $\ell_{i}\in{\bm{\ell}}$ is at least twice the differentiable loss function for the agent $i$. Differentiable losses exhibit the mixed motivation property: minimization of individual losses can result in a conflict between individuals or between individual and collective objectives (e.g., maximizing individual stats and winning the game are often conflict in basketball matches). In addition to the simultaneous gradient ${\bm{\xi}}({\bm{w}})$ of individual losses with respect to the parameters of the respective players, we use ${\bm{\xi}}_{c}({\bm{w}})$ to refer to the gradient of the collective loss: ${\bm{\xi}}_{c}({\bm{w}})=\left(\nabla_{{\bm{w}}_{1}}\ell_{c},\ldots,\nabla_{{\bm{w}}_{n}}\ell_{c}\right).$
Alignment Dilemma in DMG. Direct optimization of individual losses and collective loss are straightforward ideas to solving DMG problems. However, minimizing only the individual loss of each agent is unlikely to achieve a collective optimum (Leibo et al., 2017; McKee et al., 2020b). Conversely, optimizing the collective loss might produce better overall outcomes but risks neglecting individual interests (Roesch et al., 2024). Additionally, the local convergence of gradient descent in singular collective functions is not always guaranteed (Balduzzi et al., 2018).
Example 4.1 gives a two-player differentiable mixed-motive game. We provide a visual representation of optimization trajectories using a series of methods to investigate the learning dynamics involved in resolving the example, as shown in Fig. 1.
@@
We first implement simultaneous optimization of individual losses (Simul-Ind) with respect to each parameter, defined by the learning rule ${\bm{w}}_{i}={\bm{w}}_{i}-\gamma{\bm{\xi}}_{i}$, for $i\in\{1,2\}$, where ${\bm{\xi}}_{i}=\nabla_{{\bm{w}}_{i}}\ell_{i}$. Simultaneous optimization of collective loss (Simul-Co) replaces the individual gradient with the collective gradient ${\bm{\xi}}_{c}$ of collective loss $\ell_{c}=\ell_{1}+\ell_{2}$. Furthermore, in order to investigate the learning dynamics of prevalent gradient modification optimization approaches such as SGA and CGA for the two-player differentiable mixed-motive game, we modify the learning rule by integrating the respective adjusted gradient as delineated in Section 3.2.
@@
Figure 2: Left figure a: Illustration of Corollary 4.3. In case 1, within an unstable fixed point’s neighborhoods, an appropriate selection of the $\lambda$ sign push AgA to evade the unstable fixed point and pull towards a stable fixed point in its neighborhoods as shown in case 2. Right figure b: Alignment Effectiveness of AgA. The comparison between AgA (shown in red) and AgA without sign alignment (AgA-Sign, in purple) trajectories spans 40 steps, marked at every tenth step. Norm gradients are represented with blue arrows. Starting from the 14th step, sign alignment pulls the gradient toward the steepest direction, resulting in AgA reducing the number of steps by approximately 15% compared to AgA-Sign by the end of the trajectory.
@@ 4.2 Altruistic Gradient Adjustment
To tackle the challenge of aligning objectives in DMG, we propose the Altruistic Gradient Adjustment (AgA) method. Unlike existing gradient adjustment techniques (Mescheder et al., 2017; Balduzzi et al., 2018; Letcher et al., 2019; Chen et al., 2023), which primarily aim at achieving stable fixed points for individual objectives, our AgA method simultaneously considers both individual and collective objectives when seeking stable fixed points for the collective objective. Specifically, AgA is defined as follows.
@@
Note that the Hessian matrix ${\bm{H}}_{c}$ is symmetric. For brevity, we denote $\nabla{\mathcal{H}}_{c}$ as $\nabla_{\bm{w}}\left(\frac{1}{2}\|{\bm{\xi}}_{c}\|^{2}\right)$, which simplifies to ${\bm{H}}_{c}^{T}{\bm{\xi}}_{c}$. In our AgA method, the formula $\xi_{c}+\lambda(\xi+{\bm{H}}^{T}_{c}\xi_{c})$ includes the Hessian matrix, but we do not compute it directly. Instead, we calculate Hessian-vector products ${\bm{H}}^{T}_{c}\xi_{c}$, which suffice for determining the adjustment gradient and the sign of $\lambda$. This approach reduces computational complexity, enhancing the effectiveness of the AgA method. The cost for computing Hessian-vector products ${\bm{H}}^{T}_{c}\xi_{c}$ is $\mathcal{O}(n)$ for $n$ weights (Pearlmutter, 1994).
While AgA introduces additional complexity, we theoretically demonstrate that an appropriate choice of the sign of $\lambda$ ensures that AgA pulls the gradient towards a stable fixed point through the adjustment term in the neighborhood of fixed points. Conversely, when dealing with an unstable fixed point, AgA will push the gradient away from these unstable points.
Before introducing the corollary, we first provide some basic notations. The inner product of vectors ${\bm{a}}$ and ${\bm{b}}$ are denoted by $\langle{\bm{a}},{\bm{b}}\rangle$. If the Hessian matrix ${\bm{H}}$ is non-negative-definite, $\langle{\bm{\xi}}_{c},\nabla{\mathcal{H}}\rangle\geq 0$ for a non-zero ${\bm{\xi}}_{c}$. Analogously, if ${\bm{H}}$ is negative-definite, $\langle{\bm{\xi}}_{c},\nabla{\mathcal{H}}\rangle<0$ for a non-zero ${\bm{\xi}}_{c}$. Lastly, the angle between the two vectors ${\bm{a}}$ and ${\bm{b}}$ is denoted by $\theta({\bm{a}},{\bm{b}})$. Then, we state the corollary as follows:
@@
In the neighborhood of fixed points of the collective objective, AgA will pull the gradient toward stable fixed points, which means $\theta(\tilde{{\bm{\xi}}},\nabla{\mathcal{H}}_{c})\leq\theta({\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c})$, and push away from unstable ones, indicated by $\theta(\tilde{{\bm{\xi}}},\nabla{\mathcal{H}}_{c})\geq\theta({\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c})$, if $\lambda$ satisfies $\lambda\cdot\langle{\bm{\xi}}_{c},\nabla{\mathcal{H}}_{c}\rangle(\langle{\bm{\xi}},\nabla{\mathcal{H}}_{c}\rangle+\|\nabla{\mathcal{H}}_{c}\|^{2})\geq 0$.
@@
The proof is provided in Appendix B. Fig. 2(a) presents an illustrative visualization of Corollary 4.3. The figure depicts two scenarios: the neighborhoods of stable and unstable fixed point (denoted by a star), respectively. In the neighborhood of an unstable fixed point of the collective objective, as depicted in Case 1, our AaA gradient $\tilde{{\bm{\xi}}}$ is pushed out of the region, resulting in a faster escape compared to the original collective gradient ${\bm{\xi}}_{c}$. This behavior is exemplified by the green arrow. Conversely, Case 2 illustrates the scenario of a stable fixed point, wherein our AaA gradient $\tilde{{\bm{\xi}}}$ is pulled towards the stable equilibrium, exemplified by the red arrow.
AgA could be easily integrated into any centralized training with decentralized execution (CTDE) framework within MARL. Detailed statements on the implementation of AgA, including the pseudocode, are provided in Appendix C.
4.3 Alignment Effectiveness of AgA: A Toy Experiment
@@
To address the two-player DMG as outlined in Example 4.1, we incorporate the AgA method into the fundamental gradient descent algorithm. The optimization trajectories in both the collective reward landscape and the individual player reward contour are represented in Fig. 1(a), Fig. 1(b), and Fig. 1(c). In contrast to the objective misalignment exhibited by Simul-Co, AgA successfully aligns the agents’ interests, as evidenced in Fig. 1(b) and Fig 1(c). Fig. 1(b) depicts the trajectory of AgA meticulously carving its course along the summit of the individual reward contour. Furthermore, AgA is also shown to improve Player 2 reward (as shown in Fig. 1(c)), despite a slower convergence rate than Simul-Co.
@@
In addition, Fig. 2(b) presents a critical comparison between the trajectories of AgA (in red) and AgA without sign alignment (AgA-Sign, depicted in purple), as outlined in Corollary 4.3. This side-by-side comparison covers 40 steps and features highlighted points every tenth step. It provides a visual illustration of the effectiveness introduced by sign alignment starting from the 14th step in the trajectories of AgA and AgA without sign alignment. Remarkably, norm gradients are represented by blue arrows, indicating the direction of the fastest updates. As depicted in the figure, AgA with sign selection is aligned toward the steepest update direction, resulting in a reduction of approximately 15% in the number of steps compared to AgA-Sign by the end of the trajectory.
@@ 5 Experiments
To assess the effectiveness of the proposed AgA algorithm, we perform comprehensive experiments across various environments, ranging from simple to complex, and from small to large scales. The initial environment involves a two-player public goods game (see Section 5.1), followed by sequential social dilemma environments (see Section 5.2): Cleanup and Harvest, involving 5 homogeneous players. The final test is conducted in our specially developed selfish-MMM2 environment (see Section 5.3), which is both more complex and larger in scale, featuring approximately 10 controlled agents of 3 distinct types, competing against 11 adversaries. Comparative experiments employ commonly used baseline approaches, such as simultaneous optimization with individual and collective losses (termed Simul-Ind and Siml-Co), CGA (Mescheder et al., 2017), SGA (Balduzzi et al., 2018), SVO (McKee et al., 2020a), and the selfishness level approach (abbreviated as SL) (Roesch et al., 2024).